# Rotation sets and axes in the fine curve graph for torus homeomorphisms

### v1

@@ 10.4. Metric characterization of fixed points of loxodromic
In this section we describes examples of torus homeomorphisms that illustrate our results. In section 11.1 we give formulas for homeomorphisms having a given parallelogram as a rotation set. For triangles we are not aware of any formula, thus instead we provide a geometric construction in the spirit of Kwapisz [Kwa92], in section 11.2. In section 11.3 we define the twist numbers between two points at infinity, and the twist number of a loxodromic element, and prove some elementary properties. We illustrate this notion in section 11.4 by providing an example of computation for a nice family of examples. In section 11.5 we describe a free subgroup of $\mathrm{homeo}_{0}(T^{2})$ made of loxodromic elements, whose scaled rotation sets behave as badly as possible: some boundary points in $\partial_{\infty}\mathcal{C}^{\dagger}(T^{2})$ have no associated (scaled) rotation sets, and the scaled rotation set associated to the attracting fixed points of loxodromic element does not depend continuously on the element.
@@ 11.1. Parallelograms
We first explain how to realize the square $0,1]^{2}$ by a homeomorphism $h$ which is conjugate to its inverse, from which we will deduce the general case. Let $\tilde{f}(x,y)=(x+\sin^{2}(\pi y),y)$, $\tilde{g}(x,y)=(x,y+\sin^{2}(\pi x))$, $\tilde{h}=\tilde{g}\tilde{f}$. Then the rotation set of $\tilde{h}$ is the square $0,1]^{2}$ (see for instance [LPSW24]). Furthermore, denoting $T_{u}$ the translation by vector $u$, we note that the map $T_{(\frac{1}{2},0)}$ commutes with $\tilde{f}$ and conjugate $\tilde{g}$ to $(x,y+1-\sin^{2}(\pi x))$. Denoting $T=T_{(\frac{1}{2},\frac{1}{2})}$, we get
@@ 11.2. Triangles
To get our example we follow the construction by Kwapisz ([Kwa92]), as exposed by Béguin in [Bé07], except that we do not require $f$ to be Axiom A (we observe that Markov partitions are not needed for computing the rotation set). We include the construction since the more general construction of Kwapisz, which allows him to realize every rational polygon, is much more difficult to follow.
@@ 11.3. Twists numbers for boundary points, and for maps
Now let $(x_{i})_{i\in\mathbb{Z}}$ be a $K$-quasigeodesic path joining $\xi$ to $\xi^{\prime}$ for some $K$ (for the existence of quasigeodesics connecting boundary points, see for instance [LT25], Proposition 2.3). For every $\alpha$ outside the $r(K)$-neighborhood of this path, $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})\leq 2$; here $r(K)$ is the constant from the Bounded Geodesic Image Theorem. We now assume that $\mathrm{tw}(\xi,\xi^{\prime})=+\infty$: by definition there is a sequence $(\alpha_{i})$ such that $\mathrm{tw}_{\alpha_{i}}(\xi,\xi^{\prime})$ tends to $\infty$. According to the previous observations, such a sequence lies inside the $r(K)$ neighborhood of our path, and escape every ball. Up to extracting, such a sequence converges either to $\xi$ or $\xi^{\prime}$. There are two (overlapping) cases: (i) there are such sequences converging to $\xi$, or (ii) there are such sequences converging to $\xi^{\prime}$.
@@
The Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ is equipped with a topology (see e.g. [BH99] for a general discussion of boundaries, and [BHW24, LT25] specifically for the fine curve graph).
@@
Let us go back to our specific context. Let $(f_{m})$ be a sequence in $\mathrm{Homeo}(S)$ which converges to $f$ for the $C^{0}$ distance. In our context, the space $X=\mathcal{C}^{\dagger}(S)$ is a graph, the distance takes only integer values, and the ”simple” convergence does not hold. Instead, we have the following ”coarse simple convergence” (cf [BHM+22, Lemma 3.5]): for every $\alpha$, and every $n$ large enough, we have $d(f_{m}\alpha,f\alpha)\leq 2$. With this property, the convergence of $(\xi^{+}(f_{m}))$ to $(\xi^{+}(f))$ may be proved along the same lines as the continuity of the translation length ([BHM+22, Theorem 3.4]).
Here is a sketch of proof, details are left to the reader. According to [BHM+22, Lemma 3.8], up to replacing each map with a fixed power, we may assume the following. There is a curve $\alpha$ such that for $m$ large enough, for every $n$, and for every $i$ with $n\leq i\leq n$, the point $f_{m}^{i}\alpha$ is within $4\delta$ of the geodesic $f_{m}^{-n}\alpha,f_{m}^{n}\alpha]$. Fix $i_{0}$ so that $f^{i_{0}}\alpha$ is close to $\xi^{+}(f)$, and $m$ large enough so that the previous property holds, and $d(f_{m}^{i_{0}}\alpha,f^{i_{0}}\alpha)\leq 2$: then the set $N=\{x\in\mathcal{C}^{\dagger}(S),(x,f_{m}^{i_{0}}\alpha)_{\alpha}\geq d(\alpha,f_{m}^{i_{0}}\alpha)-100\delta\}$ is a small neighborhood of $\xi^{+}(f)$. For every $n\geq i_{0}$, the geodesic $f_{m}^{-n}\alpha,f_{m}^{n}\alpha]$ passes close to $f_{m}^{i_{0}}\alpha$, and thus the point $f_{m}^{n}\alpha$ belongs to $N$. We conclude that $\xi^{+}(f_{m})$ belongs to the closure of $N$, which is a small neighborhood of $\xi^{+}(f)$ in the $\mathcal{C}^{\dagger}(S)\cup\partial\mathcal{C}^{\dagger}(S)$. ∎
@@
Remember that the map $f\mapsto\mathrm{tl}(f)$ is also continuous ([BHM+22, Theorem 3.4]). In particular being loxodromic is an open property. By chaining the two lemmas we immediately get the following.
@@
In the next subsection we will see that there are examples with arbitrarily large twist number. Together with the semi-continuity, this implies the following non density result. Remember that $\mathrm{Homeo}_{0}(S)$ acts minimally on the boundary, see [BH25, LT25].
@@ 11.4. Computation of twist and translation length
is an invariant quasi-axis for $h$. This already tells us that $\mathrm{tl}(h)\leq 2$, as noted in [LPSW24].
@@ 11.5. Schottky groups
Examples with these properties may be constructed the following way. Let $P$ be a rational polygon containing 0. Kwapisz [Kwa92] constructs an Axiom A diffeomorphisms $f$ whose rotation set is $P$, whose periodic orbits belong to some Cantor set, except for some sinks and sources that have rotation vector 0, and 0 is also the rotation vector of some fixed point in the Cantor set. One may blow-up one of the source or sink to get a disk $D(f)$ of fixed points disjoint from the Cantor set. Now one can choose $C(f)$ to be another disc that contains the Cantor set and is disjoint from $D(f)$. The map $g$ is constructed analogously, and then conjugated to get property 3.

### v2

@@ 11. Stabilizers of points at infinity
In this section we describes examples of torus homeomorphisms that illustrate our results. In section 12.1 we give formulas for homeomorphisms having a given parallelogram as a rotation set. For triangles we are not aware of any formula, thus instead we provide a geometric construction in the spirit of Kwapisz [kwapisz_realise], in section 12.2. In section 12.3 we define the twist numbers between two points at infinity, and the twist number of a loxodromic element, and prove some elementary properties. We illustrate this notion in section 12.4 by providing an example of computation for a nice family of examples. In section 12.5 we describe a free subgroup of $\mathrm{homeo}_{0}(T^{2})$ made of loxodromic elements, whose scaled rotation sets behave as badly as possible: some boundary points in $\partial_{\infty}\mathcal{C}^{\dagger}(T^{2})$ have no associated (scaled) rotation sets, and the scaled rotation set associated to the attracting fixed points of loxodromic element does not depend continuously on the element.
@@ 12.1. Parallelograms
We first explain how to realize the square $0,1]^{2}$ by a homeomorphism $h$ which is conjugate to its inverse, from which we will deduce the general case. Let $\tilde{f}(x,y)=(x+\sin^{2}(\pi y),y)$, $\tilde{g}(x,y)=(x,y+\sin^{2}(\pi x))$, $\tilde{h}=\tilde{g}\tilde{f}$. Then the rotation set of $\tilde{h}$ is the square $0,1]^{2}$ (see for instance [LRPSW]). Furthermore, denoting $T_{u}$ the translation by vector $u$, we note that the map $T_{(\frac{1}{2},0)}$ commutes with $\tilde{f}$ and conjugate $\tilde{g}$ to $(x,y+1-\sin^{2}(\pi x))$. Denoting $T=T_{(\frac{1}{2},\frac{1}{2})}$, we get
@@ 12.2. Triangles
To get our example we follow the construction by Kwapisz ([kwapisz_realise]), as exposed by Béguin in [beguin], except that we do not require $f$ to be Axiom A (we observe that Markov partitions are not needed for computing the rotation set). We include the construction since the more general construction of Kwapisz, which allows him to realize every rational polygon, is much more difficult to follow.
@@ 12.3. Twists numbers for boundary points, and for maps
Now let $(x_{i})_{i\in\mathbb{Z}}$ be a $K$-quasigeodesic path joining $\xi$ to $\xi^{\prime}$ for some $K$ (for the existence of quasigeodesics connecting boundary points, see for instance [LongTan], Proposition 2.3). For every $\alpha$ outside the $r(K)$-neighborhood of this path, $\mathrm{tw}_{\alpha}(\xi,\xi^{\prime})\leq 2$; here $r(K)$ is the constant from the Bounded Geodesic Image Theorem. We now assume that $\mathrm{tw}(\xi,\xi^{\prime})=+\infty$: by definition there is a sequence $(\alpha_{i})$ such that $\mathrm{tw}_{\alpha_{i}}(\xi,\xi^{\prime})$ tends to $\infty$. According to the previous observations, such a sequence lies inside the $r(K)$ neighborhood of our path, and escape every ball. Up to extracting, such a sequence converges either to $\xi$ or $\xi^{\prime}$. There are two (overlapping) cases: (i) there are such sequences converging to $\xi$, or (ii) there are such sequences converging to $\xi^{\prime}$.
@@
The Gromov boundary $\partial\mathcal{C}^{\dagger}(S)$ is equipped with a topology (see e.g. [BH] for a general discussion of boundaries, and [Dagger3, LongTan] specifically for the fine curve graph).
@@
Let us go back to our specific context. Let $(f_{m})$ be a sequence in $\mathrm{Homeo}(S)$ which converges to $f$ for the $C^{0}$ distance. In our context, the space $X=\mathcal{C}^{\dagger}(S)$ is a graph, the distance takes only integer values, and the ”simple” convergence does not hold. Instead, we have the following ”coarse simple convergence” (cf [Dagger2, Lemma 3.5]): for every $\alpha$, and every $n$ large enough, we have $d(f_{m}\alpha,f\alpha)\leq 2$. With this property, the convergence of $(\xi^{+}(f_{m}))$ to $(\xi^{+}(f))$ may be proved along the same lines as the continuity of the translation length ([Dagger2, Theorem 3.4]).
Here is a sketch of proof, details are left to the reader. According to [Dagger2, Lemma 3.8], up to replacing each map with a fixed power, we may assume the following. There is a curve $\alpha$ such that for $m$ large enough, for every $n$, and for every $i$ with $n\leq i\leq n$, the point $f_{m}^{i}\alpha$ is within $4\delta$ of the geodesic $f_{m}^{-n}\alpha,f_{m}^{n}\alpha]$. Fix $i_{0}$ so that $f^{i_{0}}\alpha$ is close to $\xi^{+}(f)$, and $m$ large enough so that the previous property holds, and $d(f_{m}^{i_{0}}\alpha,f^{i_{0}}\alpha)\leq 2$: then the set $N=\{x\in\mathcal{C}^{\dagger}(S),(x,f_{m}^{i_{0}}\alpha)_{\alpha}\geq d(\alpha,f_{m}^{i_{0}}\alpha)-100\delta\}$ is a small neighborhood of $\xi^{+}(f)$. For every $n\geq i_{0}$, the geodesic $f_{m}^{-n}\alpha,f_{m}^{n}\alpha]$ passes close to $f_{m}^{i_{0}}\alpha$, and thus the point $f_{m}^{n}\alpha$ belongs to $N$. We conclude that $\xi^{+}(f_{m})$ belongs to the closure of $N$, which is a small neighborhood of $\xi^{+}(f)$ in the $\mathcal{C}^{\dagger}(S)\cup\partial\mathcal{C}^{\dagger}(S)$. ∎
@@
Remember that the map $f\mapsto\mathrm{tl}(f)$ is also continuous ([Dagger2, Theorem 3.4]). In particular being loxodromic is an open property. By chaining the two lemmas we immediately get the following.
@@
In the next subsection we will see that there are examples with arbitrarily large twist number. Together with the semi-continuity, this implies the following non density result. Remember that $\mathrm{Homeo}_{0}(S)$ acts minimally on the boundary, see [DaggerSurvey, LongTan].
@@ 12.4. Computation of twist and translation length
is an invariant quasi-axis for $h$. This already tells us that $\mathrm{tl}(h)\leq 2$, as noted in [LRPSW].
@@ 12.5. Schottky groups
Examples with these properties may be constructed the following way. Let $P$ be a rational polygon containing 0. Kwapisz [kwapisz_realise] constructs an Axiom A diffeomorphisms $f$ whose rotation set is $P$, whose periodic orbits belong to some Cantor set, except for some sinks and sources that have rotation vector 0, and 0 is also the rotation vector of some fixed point in the Cantor set. One may blow-up one of the source or sink to get a disk $D(f)$ of fixed points disjoint from the Cantor set. Now one can choose $C(f)$ to be another disc that contains the Cantor set and is disjoint from $D(f)$. The map $g$ is constructed analogously, and then conjugated to get property 3.